Imagine a new organism with 4 c subunits in the Fo portion of ATP synthase.

(a) Assuming the rest of metabolism and the ETC are the same as in known organisms, how many ATP are produced per NADH oxidized?

(b) Describe, in qualitative terms, the relationship between the amount of ATP synthesized per NADH oxidized and the number of c subunits in ATP synthase (i.e. if there are fewer subunits, what does that do to the number of ATP molecules synthesized per NADH oxidized?)
Answer #1

a. When a gamma subunit of ATP synthase undergoes a 360 degrees rotation, there is synthesis of 3 ATP. Each ATP synthase has 4 c subunit. Thus, H+/ATP ratio= 4/3=1.3.

10 electrons are transferred from NADH through the electron transport chain. 1 proton is required for transport of ADP and Pi.

Thus, P/0 ratio= ATP synthesized= 10/ (1.3+1) = 4.35

Thus, 4.35 ATP are produced per molecule of NADH.

b. When there are 12 C subunits, there H+/ATP ratio= 12/3= 4.

Hence, P/O ratio= 10/ (4+1) = 10/5= 2

When there are 10 C subunits, there H+/ATP ratio= 10/3= 3.33.

Hence, P/O ratio= 10/ (3.33+1) = 10/4.33= 2.3.

When there are 8 C subunits, there H+/ATP ratio= 8/3= 2.67.

Hence, P/O ratio= 10/ (2.67+1) = 10/3.67= 3.74

Thus, the number of ATP produced from NADH decreases with increase in number of c subunit of Fo subunit of ATP synthase. Fewer the c subunit, higher is the amount of ATP synthesized from one molecule of NADH.